OpenStudy (anonymous):

Will someone explain to me how to solve 1/2+1/4=1/x?

OpenStudy (jdoe0001):

if you multiply both sides by 4 what do you end up with?

OpenStudy (anonymous):

Do you mean find the least common multiple?

OpenStudy (jdoe0001):

well... that's another way. yes

OpenStudy (anonymous):

I have to explain how I get to the end answer, And I am stuck after cross multiplying...

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{2}+\cfrac{1}{4}=\cfrac{1}{x}\qquad {\color{brown}{ \times 4x}} \\ \quad \\ {\color{brown}{ \cancel{ 4 }x\times }}\cfrac{1}{\cancel{ 2 }}+{\color{brown}{ \cancel{ 4 }x\times }}\cfrac{1}{\cancel{ 4 }}={\color{brown}{ 4\cancel{ x }\times }}\cfrac{1}{\cancel{ x }}\implies ?\)

OpenStudy (anonymous):

You have to fine a common denominator. In this equation it would be 4x. \[1(2x)/2(2x)+1(x)/4(x)=1(4)/x(4)\]

OpenStudy (anonymous):

Does that make sense?